A football is kicked vertically upward from the ground and a student gazing out of a window sees it coming passed him at 16ft/sec. The window is 30ft above the ground.




a. How high does the ball go above the ground?




b. How long does it take to go from a height of 30ft to the highest point?

1
Expert's answer
2021-11-14T17:11:32-0500

a and b) We know that the velocity upwards is determined by


"V=V_0-gt 
\\ \implies 0 \,ft/s=16\,ft/s -(32.2\,ft/s^2)t
\\ \implies t=\frac{16\,ft/s}{32.2\,ft/s^2}=0.497\,s"


that result allow us to calculate the height by considering that the initial velocity occurs when the height reached is 30 ft. The final height is described by

"h=h_0 + V_0t -\frac{gt^2}{2}
\\ h=30\,ft+(0.497\,s)(16\,ft/s -\frac{(32.2\,ft/s^2)(0.497\,s)}{2})
\\ h=33.98\,ft"


In conclusion, the maximum height reached from the bottom is 33.98 ft while it takes 0.497 s for the ball to reach the maximum height.


Reference:


Young, H. D., Freedman, R. A., & Ford, A. L. (2006). Sears and Zemansky's university physics (Vol. 1). Pearson education.